OpenStudy (anonymous):

Four coins are tossed simultaneously. what is the probability that exactly 2 coins land on heads? make it as an %

OpenStudy (mayankdevnani):

total outcomes =2*2*2*2=16 possible outcomes = possible outcomes = TTTT TTTH TTHT TTHH THTT THTH THHT THHH HTTT HTTH HTHT HTHH HHTT HHTH HHHT HHHH

OpenStudy (anonymous):

there are \(2^4=16\) possible outcomes when you toss four coins 16 is not such a big number that we couldn't list them all but it might be easier to know that 16 is your denominator and \(_4C_2\) is your numerator do you know how to compute that number? "four choose two"

OpenStudy (mayankdevnani):

now choose that in which exactly 2 coins land on heads =5

OpenStudy (mayankdevnani):

now, \[\large \bf probability=\frac{favourable~outcomes}{total~outcomes}=\frac{5}{16}\]

OpenStudy (mayankdevnani):

as you want in percent ,multiply by 100, \[\large \bf \frac{5}{16} \times 100=31.25~percent\]

OpenStudy (mayankdevnani):

hope you understand ! @norajayy

OpenStudy (mayankdevnani):

understood ??? @norajayy

OpenStudy (anonymous):

i get it alittle but im still confused i need more practice and its not right 31.25% should i put 0.3125%

OpenStudy (mayankdevnani):

yeah... put this`

OpenStudy (mayankdevnani):

0.3125

OpenStudy (mayankdevnani):

we need probability !!